OpenStudy (anonymous):

A small smooth fixed pulley of Weight 'w' is connected by a vracket to the wall as shown a block of weight 'W' is connected to a light inextensiblr string as shown if w/W = 1/2 the magnitude of the force excerted by the bracket on the pully is?

OpenStudy (anonymous):



OpenStudy (ghazi):

first of all calculate torque exerted by weight on pulley and torque exerted by pulley on bracket ...after that resolve the components you'll get force

OpenStudy (anonymous):

torque...))

OpenStudy (anonymous):

lol....

OpenStudy (anonymous):

that is a Pulley

OpenStudy (ghazi):

total downward force = Wg+wg=(W+w)g and cos of this will be your answer

OpenStudy (anonymous):

The answer should be \[\sqrt{13}W/2\]

OpenStudy (ghazi):

force exerted by bracket on pulley is equal and opposite to that of applied by the system on bracket

OpenStudy (ghazi):